OpenStudy (anonymous):

Is 4x – 3 = 19 true, false, or open? (

OpenStudy (anonymous):

false

OpenStudy (anonymous):

how are you saying that it is false?

OpenStudy (anonymous):

open

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so what is it ?

OpenStudy (anonymous):

yes what? i didn't ask you a yes or no question.

OpenStudy (anonymous):

yes it is give me a min

OpenStudy (anonymous):

true 'cos 11/2 if u sub x in the equation it give correct answer

OpenStudy (anonymous):

5.5 is the answer

OpenStudy (anonymous):

u men 4 x

OpenStudy (anonymous):

exactly x=5.5

OpenStudy (anonymous):

so its true ?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

still false

OpenStudy (anonymous):

exactly. it is only true if x = 5.5 they didn't tell you that x = 5.5 so the equation is open depending on what x is assigned to.

OpenStudy (anonymous):

yes it is true not false try to sub u 'll see that is true

OpenStudy (anonymous):

u r right mathaddphy

OpenStudy (anonymous):

4x - 19 = 19 IS ONLY TRUE IF X=5.5!! THEY DIDN'T TELL YOU WHAT X IS THEREFORE, WHETHER IT IS TRUE OR FALSE IS OPEN DEPENDING ON X

OpenStudy (anonymous):

4x - 3 = 19 sorry

OpenStudy (anonymous):

i get ur point

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I'm talking to 142343 who clearly is incorrect

jhonyy9 (jhonyy9):

4x=22 x=22/4 x=11/2

OpenStudy (anonymous):

i'm confused